OpenStudy (anonymous):

The perimeter of a trapezoid is 39a-7. Three sides have the following lengths: 9a, 5a +1, and 17a -6. What is the length of the fourth side?

OpenStudy (jdoe0001):

\(\bf \\textit{let's say our 4th side is }\quad n\\ perimeter = 39a-7\qquad sides\quad 9a,\quad 5a+1,\quad 17a-6,\quad n\\ \quad \\ \textit{recall that }\quad perimeter = \textit{sum of all sides}\\ \quad \\ 39a-7= 9a+(5a+1)+(17a-6)+n\\ \quad \\\implies 39a-7= 9a+5a+1+17a-6+n\) solve for "n", the 4th side will be in terms of "a"

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\bf \textit{let's say our 4th side is }\quad n\\ perimeter = 39a-7\qquad sides\quad 9a,\quad 5a+1,\quad 17a-6,\quad n\\ \quad \\ \textit{recall that }\quad perimeter = \textit{sum of all sides}\\ \quad \\ 39a-7= 9a+(5a+1)+(17a-6)+n\\ \quad \\\implies 39a-7= 9a+5a+1+17a-6+n\)

OpenStudy (jdoe0001):

solve for "n", the 4th side will be in terms of "a"